OpenStudy (anonymous):

Solve each system using substitution. y= 3x +11 y= -2x +1

OpenStudy (anonymous):

This means the same thing as y = 3x + 11 AND y = -2x + 1

myininaya (myininaya):

hey! :) 3x+11=-2x+1 can you solve this?

OpenStudy (anonymous):

So you can set them equal, as myininaya has done.

OpenStudy (anonymous):

yea

myininaya (myininaya):

:)

OpenStudy (anonymous):

so all i havta to is solve that euation?

myininaya (myininaya):

basically

myininaya (myininaya):

solving that equation will give us x

myininaya (myininaya):

and then when we get x we plug it into one of the equations given to find y

myininaya (myininaya):

therefore finding the point of intersection (x,y)

myininaya (myininaya):

so if you want to try to solve for x i will do the same and i can tell you if you are right but if you run into trouble we can solve together okay?

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

3x+11=-2x+1 +2x +2x 5x+11=1 -11 -11 5x= 12 ___ ____ 5 5

myininaya (myininaya):

so 1-11 is 12 ?

myininaya (myininaya):

what is 11-1?

OpenStudy (anonymous):

oh sryy

OpenStudy (anonymous):

-10

myininaya (myininaya):

perfect! :)

myininaya (myininaya):

so we have 5x=-10

OpenStudy (anonymous):

x= -2

myininaya (myininaya):

brilliant! :) now plug this into one of the equations given to find y like this y=3(-2)+11

myininaya (myininaya):

i just replaced x with -2 since x=-2

myininaya (myininaya):

now you are to simplify 3(-2)+11 to find y

OpenStudy (anonymous):

y=6 ?

myininaya (myininaya):

well 3(-2)+11 is y i just mean simplify lol

myininaya (myininaya):

-6+11=11-6=5 right?

OpenStudy (anonymous):

right,

OpenStudy (anonymous):

so y=5

OpenStudy (anonymous):

(-2,5) correct?

myininaya (myininaya):

that is totally correct

OpenStudy (anonymous):

Thank you!! XD

myininaya (myininaya):

if you had used the other equation you still would have got same thing by the way

myininaya (myininaya):

great job! you get a medal!

OpenStudy (anonymous):

Your the best!, thanks :D

myininaya (myininaya):

I know.

myininaya (myininaya):

Thanks. lol